# Scalable Multi-Agent Offline Reinforcement Learning and the Role of Information

### v1

@@ Abstract
Offline Reinforcement Learning (RL) focuses on learning policies solely from a batch of previously collected data. offering the potential to leverage such datasets effectively without the need for costly or risky active exploration. While recent advances in Offline Multi-Agent RL (MARL) have shown promise, most existing methods either rely on large datasets jointly collected by all agents or agent-specific datasets collected independently. The former approach ensures strong performance but raises scalability concerns, while the latter emphasizes scalability at the expense of performance guarantees. In this work, we propose a novel scalable routine for both dataset collection and offline learning. Agents first collect diverse datasets coherently with a pre-specified information-sharing network and subsequently learn coherent localized policies without requiring either full observability or falling back to complete decentralization. We theoretically demonstrate that this structured approach allows a multi-agent extension of the seminal Fitted Q-Iteration (FQI) algorithm [7] to globally converge, in high probability, to $\epsilon$-optimal policies. The convergence is subject to error terms that depend on the informativeness of the shared information. Furthermore, we show how this approach allows to bound the inherent error of the supervised-learning phase of FQI with the mutual information between shared and unshared information. Our algorithm, SCAlable Multi-agent FQI (SCAM-FQI), is then evaluated on a distributed decision-making problem. The empirical results align with our theoretical findings, supporting the effectiveness of SCAM-FQI in achieving a balance between scalability and policy performance.
@@ 1 Introduction
Offline Reinforcement Learning (RL) was shown to be a promising way towards deploying RL in real-world scenarios where interaction with the environment is prohibitive, costly, or risky [9]. For the many practical scenarios that involve multiple agents, recent results have shown that online RL algorithms can be scaled to multi-agent scenarios in a relatively straightforward way [8]. Surprisingly, still little is known about how single-agent offline algorithms behave in the presence of multi-agency, and even less is known about the guarantees that are transferred. All existing works either assume full observability, i.e., each agent should have access to the states of the whole team, or partial observability, i.e., each agent should have access to her own states only. While the former setting enforces good guarantees at the cost of several practical limitations [2], the latter one fails to provide any relevant performance guarantee [6]. To address these issues, we first propose a novel and scalable offline routine specifically addressing the multi-agent nature of the task, aiming at maintaining reasonable guarantees while offering a more scalable algorithm. We then propose an adaptation of Fitted Q-Iteration (FQI) algorithm [7], namely SCAlable Multi-agent FQI (SCAM-FQI), the first offline MARL algorithmic instantiation that is scalable both in training and in deployment, by requiring the agents to collect data-sets jointly with and by conditioning learned policies on the states of a restricted number of agents. We describe how the proposed algorithm enjoys nice convergence guarantees that either explicitly or implicitly depend on the quality of the information shared between the agents. Finally, we couple these statements with some empirical corroboration over an (almost) real-case decision-making scenario consisting of an industrial production scheduling problem.
Notation.
@@ 2 Problem Setting
In this work, we focus on Markov Games (MGs) [5], which consist of a tuple =$\mathcal{M}=\{\mathcal{N},\mathcal{S},\mathcal{A},\mathbb{P},r,\gamma,\mu\}$ with $\mathcal{N}$ being the set of agents interacting in the game, $\mathcal{S}$ and $\mathcal{A}$ be the joint state and action spaces respectively. $\mathbb{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the (joint) transition model governing the game dynamics, while $r$ is the reward function assumed to be sum-decomposable over the agents, namely $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,R^{\text{max}}],r(s,a):=\sum_{i\in\mathcal{N}}r_{i}(s,a)$. The state and action spaces are assumed to be factored over agents, namely $\mathcal{S}=\times_{i\in\mathcal{N}}\mathcal{S}_{i}$ and $\mathcal{A}=\times_{i\in\mathcal{N}}\mathcal{A}_{i}$. Finally, $\gamma$ and $\mu$ are the discount factor employed in the evaluation metric and the initial state distribution, respectively. Moreover, we will consider the existence of an information-sharing $\mathcal{G}=(\mathcal{N},\mathcal{E})$. The vertex set $\mathcal{N}=\{1,2,...,|\mathcal{N}|\}$ denotes the set of agents and the edge set $\mathcal{E}$ prescribes the communication links among agents: we define $\mathcal{N}_{i}=\{j\in\mathcal{N}|(i,j)\in\mathcal{E}\}$ as the set of agents that share their state-space information with agent $i$. Thus, we denote the set of states of agents in $\mathcal{N}_{i}$ with $\mathcal{S}_{\mathcal{N}_{i}}$ and the global policy in $\pi(a|s)=\prod_{i\in\mathcal{N}}\pi_{i}(a_{i}|s_{\mathcal{N}_{i}})$, i.e., given global state s, each agent acts independently according to its own policy $\pi_{i}:\mathcal{S}_{\mathcal{N}_{i}}\rightarrow\Delta(\mathcal{A}_{i})$. As a consequence, let us denote as $S$ the random variables corresponding to the joint state, then $S$ is distributed as $d^{\pi}_{\mu}\in\Delta_{\mathcal{S}}$, where $d^{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}Pr(s_{t}=s|\pi,\mu)$. Additionally, the graph $\mathcal{G}$ induces for each agent the so-called local transitions and rewards, namely $\mathbb{P}^{i}(s_{\mathcal{N}_{i}}^{\prime}|s_{\mathcal{N}_{i}},a_{i}):\mathcal{S}_{\mathcal{N}_{i}}\times\mathcal{A}_{i}\rightarrow\Delta(S_{\mathcal{N}_{i}})$ and $\bar{r}_{i}(s_{\mathcal{N}_{i}},a_{i}):\mathcal{S}_{\mathcal{N}_{i}}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$. As a result, we can define the decentralized Q-functions as:
@@
That will be compared against their centralized counterparts:
@@
$\bar{\mathcal{T}}^{i}f_{i}(s_{\mathcal{N}_{i}},a_{i})$ $=\bar{r}_{i}(s_{\mathcal{N}_{i}},a_{i})+\gamma\operatorname*{\mathbb{E}}_{s_{\mathcal{N}_{i}}^{\prime}\sim\mathbb{P}^{i}(\cdot|s_{\mathcal{N}_{i}},a_{i})}\max_{a_{i}^{\prime}}f_{i}(s_{\mathcal{N}_{i}}^{\prime},a_{i}^{\prime}),\quad\bar{\mathcal{T}}f(s,a):=\sum_{i\in\mathcal{N}}\bar{\mathcal{T}}^{i}f_{i}(s_{\mathcal{N}_{i}},a_{i})$ (3)
$\mathcal{T}f(s,a)$ $=r(s,a)+\gamma\operatorname*{\mathbb{E}}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\max_{a^{\prime}}f(s^{\prime},a^{\prime})$ (4)
The learning objective will be defined from the centralized point of view, that is as learning the optimal policy $\pi^{\star}(\cdot|s)\in\operatorname*{arg\,max}_{\pi}\operatorname*{\mathbb{E}}_{a\sim\pi}Q^{\star}(s,a)$ with $Q^{\star}$ being the fixed point of the centralized optimal Bellman Operator of Eq. (4).
3 Scalable Offline MARL
Offline Collection:
In our scalable multi-agent offline scenario, each agent will have access to different data-set $\mathcal{D}_{i}=\{(s_{\mathcal{N}_{i},t},a_{i,t},\bar{r}_{i,t},s^{\prime}_{\mathcal{N}_{i},t})\}$; in other words, agents will only have to collect their own experience and share it according to $\mathcal{G}$, without any joint collection happening. For the sake of the theoretical analysis, we will then assume that there exists a joint dataset $\mathcal{D}=\{(s_{t},a_{t},r_{t},s^{\prime}_{t})\}$ collected according to $(s,a)\sim\nu\in\Delta(\mathcal{S}\times\mathcal{A})$, $r=r(s,a)$, $s^{\prime}\sim\mathbb{P}(\cdot|s,a)$ and that it enjoys a Concentrability Assumption: $\forall\pi,\forall(s,a),\;\exists\;C:\;\frac{d^{\pi}(s,a)}{\nu(s,a)}\leq C,$. One should note that the actual size of the per-agent datasets will strictly depend on $\mathcal{G}$, and we will have, in general, that $|\mathcal{D}|<<|\mathcal{D}_{i}|$.
@@ 3 Scalable Offline MARL
The proposed algorithm SCAM-FQI consists of a scalable multi-agent instantiation of FQI [7]; it performs the following simple iteration: starting with some function $\hat{q}^{0}_{i}\in\mathcal{F}_{i}$ over $\mathcal{F}_{i}=\{f:\mathcal{S}_{\mathcal{N}_{i}}\times\mathcal{A}_{i}\rightarrow[0,V^{\text{max}}]\},V^{\text{max}}=\frac{R^{\text{max}}}{1-\gamma}$, each agent at iteration $k$ independently performs the following updates:
@@
Convergence Guarantees:
In the following, we show with a proof following standard techniques [1] that SCAM-FQI leverages the information-sharing graph so as to maintain good performance guarantees. First of all:
@@
We start by considering the Performance Difference Lemma between joint functions against the optimal policy $\pi^{\star}$ [^1] Here we simplified the notation by denoting $A^{\star}(s,\pi^{k}(s)):=\operatorname*{\mathbb{E}}_{a\sim\pi^{k}(\cdot|s)}A^{\star}(s,a)$ and dropping the dependency over the fixed initial distribution $\mu$.:
$(1-\gamma)(V^{\star}_{\mu}-V^{\pi^{k}}_{\mu})$ $=\operatorname*{\mathbb{E}}_{s\sim d^{\pi^{k}}}[Q^{\star}(s,\pi^{\star}(s))-Q^{\star}(s,\pi^{k}(s))]$
$\leq\operatorname*{\mathbb{E}}_{s\sim d^{\pi^{k}}}[Q^{\star}(s,\pi^{\star}(s))-\hat{q}^{\pi^{k}}(s,\pi^{\star}(s))+\hat{q}^{\pi^{k}}(s,\pi^{k}(s))-Q^{\star}(s,\pi^{k}(s))]$
@@
Now, for a generic term $\left\|Q^{\star}-\hat{q}^{\pi^{k}}\right\|_{1,\beta}\leq\left\|Q^{\star}-{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}+\left\|\hat{q}^{\pi^{k}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}+\left\|\mathcal{T}\hat{q}^{\pi^{k-1}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}$, we have that:
$\left\|Q^{\star}-{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}$ $\leq\gamma\operatorname*{\mathbb{E}}_{(s,a)\sim\beta}\big{|}\operatorname*{\mathbb{E}}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\max_{a^{\prime}}Q^{\star}(s^{\prime},a^{\prime})-\max_{a^{\prime}}\hat{q}^{\pi^{k-1}}(s^{\prime},a^{\prime})\big{|}$
$\leq\gamma\left\|Q^{\star}-\hat{q}^{\pi^{k-1}}\right\|_{1,\beta^{\prime\prime}},\beta^{\prime\prime}(s^{\prime},a^{\prime})=\sum_{s,a}\beta(s,a)\mathbb{P}(s^{\prime}|s,a)1\{a^{\prime}=\operatorname*{arg\,max}_{a}|Q^{\star}(s^{\prime},a)-\hat{q}^{\pi^{k-1}}(s^{\prime},a)|\}$
$\left\|\mathcal{T}\hat{q}^{\pi^{k-1}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}$ $=\operatorname*{\mathbb{E}}_{(s,a)\sim\beta}\big{|}\sum_{i}r_{i}(s,a)+\gamma\operatorname*{\mathbb{E}}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\max_{a^{\prime}}\hat{q}^{\pi^{k-1}}(s^{\prime},a^{\prime})-\sum_{i}\bar{r}_{i}(s_{\mathcal{N}_{i}},a_{i})-\gamma\operatorname*{\mathbb{E}}_{s^{\prime}_{\mathcal{N}_{i}}\sim\mathbb{P}^{i}(\cdot|s_{\mathcal{N}_{i}},a_{i})}\max_{a_{i}^{\prime}}\hat{q}^{\pi^{k-1}_{i}}_{i}(s^{\prime}_{\mathcal{N}_{i}},a_{i}^{\prime}))\big{|}$
@@
$\left\|\hat{q}^{\pi^{k}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}$ $\leq\left\|\hat{q}^{\pi^{k}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{2,\beta}\leq\sum_{i}\sqrt{C}\left\|\hat{q}^{\pi_{i}^{k}}-\bar{\mathcal{T}}^{i}\hat{q}^{\pi_{i}^{k-1}}\right\|_{2,\nu}$
@@
Fortunately enough, even assuming to be performing simple least-squares estimation in the Q-fitting step, standard generalization bounds hold for the $\epsilon$ error term of Lemma 3.1:
Lemma 3.2 (Least Squares Generalization Bound [4]).
@@
Define the value functions $V^{\pi}_{\mu}:=\operatorname*{\mathbb{E}}_{s\sim d^{\pi}_{\mu},a\sim\pi(\cdot|s)}Q^{\pi}(s,a)$, assume that $\forall i,j\in\mathcal{N}\mathcal{F}_{i}=\mathcal{F}_{j}=\mathcal{F},i\neq j$, fix $K\in\mathbb{N}^{+}$. The SCAM-FQI Iteration of Eqs. (5) guarantees that with probability $1-\delta$ for any initial distribution $\mu\in\Delta(\mathcal{S})$ it holds
$V^{\star}_{\mu}-V^{\pi^{K}}_{\mu}\leq\frac{2\gamma^{K-1}}{1-\gamma}(\gamma V^{\text{max}}+\sqrt{C}\epsilon_{r}+\frac{\sqrt{C}\gamma}{1-\gamma}\epsilon_{\mathbb{P}})+\frac{2N}{(1-\gamma)^{2}}\Bigg{(}\sqrt{\frac{22C(V^{\text{max}})^{2}\ln(|\mathcal{F}|KN/\delta)}{|D_{i}|}}+\sqrt{20\epsilon_{\text{INH}}}\Bigg{)}.$ (8)
In other words, we are able to almost fully recover the convergence guarantees attained in the single-agent cases out of some core differences: $(i)$ the two error terms $\epsilon_{r}$ and $\epsilon_{\mathbb{P}}$ represent the bias induced by the per-agent local transitions and reward functions, and they are indeed controlled by the ability of the network $\mathcal{G}$ to recover the true transition and reward models; $(ii)$ the second term apparently scales linearly with the number of agents, yet one should notice that the nature of the communication network indirectly induces a trade-off on it as well, as the size of the actual per-agent dataset $|\mathcal{D}_{i}|$ will increase as the set $\mathcal{N}_{i}$ decreases; $(iii)$ one more refined analysis of the last supervised-learning related term reveals a trade-off on the information-sharing network as well:

### v2

@@ Abstract
Offline Reinforcement Learning (RL) focuses on learning policies solely from a batch of previously collected data. offering the potential to leverage such datasets effectively without the need for costly or risky active exploration. While recent advances in Offline Multi-Agent RL (MARL) have shown promise, most existing methods either rely on large datasets jointly collected by all agents or agent-specific datasets collected independently. The former approach ensures strong performance but raises scalability concerns, while the latter emphasizes scalability at the expense of performance guarantees. In this work, we propose a novel scalable routine for both dataset collection and offline learning. Agents first collect diverse datasets coherently with a pre-specified information-sharing network and subsequently learn coherent localized policies without requiring either full observability or falling back to complete decentralization. We theoretically demonstrate that this structured approach allows a multi-agent extension of the seminal Fitted Q-Iteration (FQI) algorithm (riedmilled2005fittedqiteration, ) to globally converge, in high probability, to $\epsilon$-optimal policies. The convergence is subject to error terms that depend on the informativeness of the shared information. Furthermore, we show how this approach allows to bound the inherent error of the supervised-learning phase of FQI with the mutual information between shared and unshared information. Our algorithm, SCAlable Multi-agent FQI (SCAM-FQI), is then evaluated on a distributed decision-making problem. The empirical results align with our theoretical findings, supporting the effectiveness of SCAM-FQI in achieving a balance between scalability and policy performance.
@@ 1 Introduction
Offline Reinforcement Learning (RL) was shown to be a promising way towards deploying RL in real-world scenarios where interaction with the environment is prohibitive, costly, or risky mopo . For the many practical scenarios that involve multiple agents, recent results have shown that online RL algorithms can be scaled to multi-agent scenarios in a relatively straightforward way thesurprisingeffectiveness . Surprisingly, little is known about how single-agent offline algorithms behave in the presence of multiple agents, and even less is known about which theoretical guarantees do transfer. Existing works either assume full observability, i.e., each agent accesses the states of the whole team, or partial observability, i.e., each agent accesses her own states only. While the former setting enforces good guarantees at the cost of several practical limitations bardemodelbased , the latter one fails to provide any relevant performance guarantee Pan2021PlanBA . To address these issues, we first propose a novel and scalable offline routine in which agents collect distinct datasets, yet they partially share their respective states according to a communication graph. We then propose a multi-agent adaptation of Fitted Q-Iteration (FQI) algorithm (riedmilled2005fittedqiteration, ), namely SCAlable Multi-agent FQI (SCAM-FQI), the first offline MARL algorithmic instantiation that is scalable both in training and in deployment, as both value functions and policies are supported or conditioned over the states of a restricted number of agents according to the communication graph. We describe how the proposed algorithm enjoys nice convergence guarantees that either explicitly or implicitly depend on the quality of the information shared between the agents. Finally, we couple these statements with some empirical corroboration over an (almost) real-case decision-making scenario consisting of an industrial production scheduling problem.
@@ 2 Problem Setting
In this work, we focus on Markov Games (MGs, Littman1994, ), which consist of a tuple =$\mathcal{M}=\{\mathcal{N},\mathcal{S},\mathcal{A},\mathbb{P},r,\gamma,\mu\}$ with $\mathcal{N}$ being the set of agents interacting in the game, $\mathcal{S}$ and $\mathcal{A}$ be the joint state and action spaces respectively. $\mathbb{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the (joint) transition model governing the game dynamics, while $r$ is the reward function assumed to be sum-decomposable over the agents, namely $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,R^{\text{max}}],r(s,a):=\sum_{i\in\mathcal{N}}r_{i}(s,a)$. The state and action spaces are assumed to be factored over agents, namely $\mathcal{S}=\times_{i\in\mathcal{N}}\mathcal{S}_{i}$ and $\mathcal{A}=\times_{i\in\mathcal{N}}\mathcal{A}_{i}$. Finally, $\gamma$ and $\mu$ are the discount factor employed in the evaluation metric and the initial state distribution, respectively. We will consider the existence of an information-sharing graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$. The vertex set $\mathcal{N}=\{1,2,...,|\mathcal{N}|\}$ denotes the set of agents and the edge set $\mathcal{E}$ prescribes the communication links among agents: we define $\mathcal{N}_{i}=\{j\in\mathcal{N}|(i,j)\in\mathcal{E}\}$ as the set of agents that share their state-space information with agent $i$. Thus, we denote the set of states of agents in $\mathcal{N}_{i}$ with $\mathcal{S}_{\mathcal{N}_{i}}$ and the global policy is $\pi(a|s)=\prod_{i\in\mathcal{N}}\pi_{i}(a_{i}|s_{\mathcal{N}_{i}})$, i.e., given global state s, each agent acts independently according to its own policy $\pi_{i}:\mathcal{S}_{\mathcal{N}_{i}}\rightarrow\Delta(\mathcal{A}_{i})$. Additionally, the graph $\mathcal{G}$ induces for each agent the so-called local transitions and rewards, namely $\mathbb{P}^{i}(s_{\mathcal{N}_{i}}^{\prime}|s_{\mathcal{N}_{i}},a_{i}):\mathcal{S}_{\mathcal{N}_{i}}\times\mathcal{A}_{i}\rightarrow\Delta(S_{\mathcal{N}_{i}})$ and $\bar{r}_{i}(s_{\mathcal{N}_{i}},a_{i}):\mathcal{S}_{\mathcal{N}_{i}}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$. Finally, joint states are distributed according to $d^{\pi}_{\mu}\in\Delta_{\mathcal{S}}$, where $d^{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}Pr(s_{t}=s|\pi,\mu)$. As a result, we can define the decentralized Q-functions as:
@@
These will be compared against their centralized counterparts:
@@
$\bar{\mathcal{T}}^{i}f_{i}(s_{\mathcal{N}_{i}},a_{i})=\bar{r}_{i}(s_{\mathcal{N}_{i}},a_{i})+\gamma\operatorname*{\mathbb{E}}_{s_{\mathcal{N}_{i}}^{\prime}\sim\mathbb{P}^{i}(\cdot|s_{\mathcal{N}_{i}},a_{i})}\max_{a_{i}^{\prime}}f_{i}(s_{\mathcal{N}_{i}}^{\prime},a_{i}^{\prime}),$ $\mathcal{T}f(s,a)=r(s,a)+\gamma\operatorname*{\mathbb{E}}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\max_{a^{\prime}}f(s^{\prime},a^{\prime})$ (3)
where $\bar{\mathcal{T}}f(s,a):=\sum_{i\in\mathcal{N}}\bar{\mathcal{T}}^{i}f_{i}(s_{\mathcal{N}_{i}},a_{i})$. The objective will be defined in centralized terms, that is as learning the optimal policy $\pi^{\star}(\cdot|s)\in\operatorname*{arg\,max}_{\pi}\operatorname*{\mathbb{E}}_{a\sim\pi}Q^{\star}(s,a)$ with $Q^{\star}$ being the fixed point of the centralized optimal Bellman Operator.
3 Towards Scalable Offline MARL
In the following, we propose an offline learning approach that integrates a decentralized architecture with a structured communication network. Each agent collects a distinct dataset comprising its own local information and limited (but potentially relevant) information from a subset of other agents, and subsequently learns a policy conditioned only on this subset. Intuitively, restricting the information accessible to each agent introduces a potential performance trade-off. We will show that the degradation in performance is proportional to the extent to which the local transitions and reward approximate the true underlying dynamics. Crucially, omitting information that is either weakly relevant or highly redundant, quantified via Conditional Mutual Information (CMI), does not significantly impact the overall error. This selective conditioning also enables the exploitation of substantially larger per-agent datasets, thus improving sample efficiency without compromising performance guarantees.
Data Collection:
Each agent will have access to different data-set $\mathcal{D}_{i}=\{(s_{\mathcal{N}_{i},t},a_{i,t},\bar{r}_{i,t},s^{\prime}_{\mathcal{N}_{i},t})\}$; in other words, agents will only have to collect their own experience and share it according to $\mathcal{G}$, without any joint collection happening. For the sake of the theoretical analysis, we will then assume that there exists a joint dataset $\mathcal{D}=\{(s_{t},a_{t},r_{t},s^{\prime}_{t})\}$ collected according to $(s,a)\sim\nu\in\Delta(\mathcal{S}\times\mathcal{A})$, $r=r(s,a)$, $s^{\prime}\sim\mathbb{P}(\cdot|s,a)$ and that it enjoys a Concentrability Assumption: $\forall\pi,\forall(s,a),\;\exists\;C:\;\frac{d^{\pi}(s,a)}{\nu(s,a)}\leq C$. Crucially, the actual size of the datasets will depend on $\mathcal{G}$, and that, in general, $|\mathcal{D}|<<|\mathcal{D}_{i}|$.
@@ 3 Towards Scalable Offline MARL
The proposed algorithm SCAM-FQI consists of a scalable multi-agent instantiation of FQI (riedmilled2005fittedqiteration, ); it performs the following simple iteration: starting with some function $\hat{q}^{0}_{i}\in\mathcal{F}_{i}$ over $\mathcal{F}_{i}=\{f:\mathcal{S}_{\mathcal{N}_{i}}\times\mathcal{A}_{i}\rightarrow[0,V^{\text{max}}]\},V^{\text{max}}=\frac{R^{\text{max}}}{1-\gamma}$, each agent at iteration $k$ independently performs the following updates:
@@
Convergence Guarantees:  Now, we show with a proof following standard techniques (Agarwal2019ReinforcementLT, ) that SCAM-FQI leverages the information-sharing graph so as to maintain good performance guarantees. First of all:
@@
We start with the Performance Difference Lemma between joint functions against the optimal policy $\pi^{\star}$ [^1] Here we simplified the notation by denoting $A^{\star}(s,\pi^{k}(s)):=\operatorname*{\mathbb{E}}_{a\sim\pi^{k}(\cdot|s)}A^{\star}(s,a)$ and dropping the dependency over the fixed initial distribution $\mu$.:
$(1-\gamma)(V^{\star}_{\mu}-V^{\pi^{k}}_{\mu})$ $=\operatorname*{\mathbb{E}}_{s\sim d^{\pi^{k}}}[Q^{\star}(s,\pi^{\star}(s))-Q^{\star}(s,\pi^{k}(s))]\leq\operatorname*{\mathbb{E}}_{s\sim d^{\pi^{k}}}[Q^{\star}(s,\pi^{\star}(s))-\hat{q}^{\pi^{k}}(s,\pi^{\star}(s))+\hat{q}^{\pi^{k}}(s,\pi^{k}(s))-Q^{\star}(s,\pi^{k}(s))]$
@@
Now, for a generic term we have that $\left\|Q^{\star}-\hat{q}^{\pi^{k}}\right\|_{1,\beta}\leq\overset{(a)}{\left\|Q^{\star}-{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}}+\overset{(b)}{\left\|\hat{q}^{\pi^{k}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}}+\overset{(c)}{\left\|\mathcal{T}\hat{q}^{\pi^{k-1}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}}$, and
$\left\|Q^{\star}-{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}$ $\overset{(a)}{\leq}\gamma\operatorname*{\mathbb{E}}_{(s,a)\sim\beta}\big{|}\operatorname*{\mathbb{E}}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\max_{a^{\prime}}Q^{\star}(s^{\prime},a^{\prime})-\max_{a^{\prime}}\hat{q}^{\pi^{k-1}}(s^{\prime},a^{\prime})\big{|}$
$\leq\gamma\left\|Q^{\star}-\hat{q}^{\pi^{k-1}}\right\|_{1,\beta^{\prime\prime}},\beta^{\prime\prime}(s^{\prime},a^{\prime})=\sum_{s,a}\beta(s,a)\mathbb{P}(s^{\prime}|s,a)\mathbf{1}\{a^{\prime}=\operatorname*{arg\,max}_{a}|Q^{\star}(s^{\prime},a)-\hat{q}^{\pi^{k-1}}(s^{\prime},a)|\}$
$\left\|\mathcal{T}\hat{q}^{\pi^{k-1}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}$ =$\overset{(b)}{=}\operatorname*{\mathbb{E}}_{(s,a)\sim\beta}\big{|}\sum_{i}r_{i}(s,a)+\gamma\operatorname*{\mathbb{E}}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}\max_{a^{\prime}}\hat{q}^{\pi^{k-1}}(s^{\prime},a^{\prime})-\sum_{i}\bar{r}_{i}(s_{\mathcal{N}_{i}},a_{i})-\gamma\operatorname*{\mathbb{E}}_{s^{\prime}_{\mathcal{N}_{i}}\sim\mathbb{P}^{i}(\cdot|s_{\mathcal{N}_{i}},a_{i})}\max_{a_{i}^{\prime}}\hat{q}^{\pi^{k-1}_{i}}_{i}(s^{\prime}_{\mathcal{N}_{i}},a_{i}^{\prime}))\big{|}$
@@
$\left\|\hat{q}^{\pi^{k}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{1,\beta}$ $\overset{(c)}{\leq}\left\|\hat{q}^{\pi^{k}}-\bar{\mathcal{T}}\hat{q}^{\pi^{k-1}}\right\|_{2,\beta}\leq\sum_{i}\sqrt{C}\left\|\hat{q}^{\pi_{i}^{k}}-\bar{\mathcal{T}}^{i}\hat{q}^{\pi_{i}^{k-1}}\right\|_{2,\nu}$
@@
Now, even assuming simple least-squares estimation, generalization bounds hold for the $\epsilon$ error term of Lem. 3.1:
Lemma 3.2 (Least Squares Generalization Bound kakaderandomdesign ).
@@
Define the value functions $V^{\pi}_{\mu}:=\operatorname*{\mathbb{E}}_{s\sim d^{\pi}_{\mu},a\sim\pi(\cdot|s)}Q^{\pi}(s,a)$, assume that $\forall i,j\in\mathcal{N}\mathcal{F}_{i}=\mathcal{F}_{j}=\mathcal{F},D_{i}=D_{j}=D,i\neq j$, fix $K\in\mathbb{N}^{+}$. The SCAM-FQI Iteration of Eqs. (4) guarantees that with probability $1-\delta$ for any initial distribution $\mu\in\Delta(\mathcal{S})$ it holds
$V^{\star}_{\mu}-V^{\pi^{K}}_{\mu}\leq\frac{2\gamma^{K-1}}{1-\gamma}(\gamma V^{\text{max}}+\sqrt{C}\epsilon_{r}+\frac{\sqrt{C}\gamma}{1-\gamma}\epsilon_{\mathbb{P}})+\frac{2N}{(1-\gamma)^{2}}\Bigg{(}\sqrt{\frac{22C(V^{\text{max}})^{2}\ln(|\mathcal{F}|KN/\delta)}{|D|}}+\sqrt{20\epsilon_{\text{INH}}}\Bigg{)}.$ (7)
In other words, we are able to almost fully recover the convergence guarantees attained in the single-agent cases out of some core differences: $(i)$ the two error terms $\epsilon_{r}$ and $\epsilon_{\mathbb{P}}$ represent the bias induced by the per-agent local transitions and reward functions, and they are indeed controlled by the ability of the network $\mathcal{G}$ to recover the true transition and reward models; $(ii)$ the second term apparently scales linearly with the number of agents, yet one should notice that the nature of the communication network indirectly induces a trade-off on it, as the size of the actual per-agent dataset $|\mathcal{D}_{i}|$ will increase as the set $\mathcal{N}_{i}$ decreases; $(iii)$ one more refined analysis of the last supervised-learning related term reveals a trade-off on the information-sharing network as well: